Problem 4 eld and gains 1.54 x iƠ15 J of electrostatic potential energy. Calculate the magnitude of the electrostatic potential difference An electron is moved rom point A to pointB in a uniform electric between the two points Due Friday February 01 11:59 am (EST) Submit Answer Tries 0/6HELP PLS!!
Answer #1

here,

magnitude of charge on electron , q = 1.6 * 10^-19 C

let the potential difference between the points be V

the electrostatic energy gained , U = q * V

1.54 * 10^-4 = 1.6 * 10^-19 * V

solving for V

V = 9625 V

the potential difference between the points is 9625 V